How many gallons of a 25 percent saline solution and a 35 percent solution must be mixed together to get 220 Gallons of 30 percent saline solution.
If you could send me the equation you plunged the info into as well it would be greatly appreciated. Thanks

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

How many gallons of a 25 percent saline solution and a 35 percent solution must be mixed together to get 220 Gallons of 30 percent saline solution.
-----
Equation:
salt + salt = salt
0.25x + 0.35(220-x) = 0.30*220
------------
25x + 35*220 - 35x = 30*220
-----------
-10x = -5*220
---
x = (1/2)220
---
x = 110 Gallons (amt. of 25% solution needed)
220-x = 110 Gallons (amt. of 35% solution needed)
-------------
Cheers,
Stan H.
----------

Answer by ikleyn(52776)

.
See also the lesson  Solving typical word problems on mixtures for solutions  in this site where the similar problem was solved.